Answer the following questions for the function
f(x)=x^3/(x^2−4)
defined on the interval [−19,16].

a.) Enter the x-coordinates of the vertical asymptotes of f(x) as a comma-separated list. That is, if there is just one value, give it; if there are more than one, enter them separated commas; and if there are none, enter NONE .

b)f(x) is concave up on the region=
1
Expert's answer
2020-11-24T17:35:17-0500

a) Since "\lim_{x\to\pm 2}\frac{x^3}{x^2-4}=\infty", the lines "x=-2" and "x=2" are the vertical asymptotes of "f(x)".


Answer: -2, 2


Let us find fist and second derivatives of the function "f(x)=\frac{x^3}{x^2-4}" defined on the interval "[\u221219,16]" :


"f'(x)=\frac{3x^2(x^2-4)-x^32x}{(x^2-4)^2}=\frac{x^4-12x^2}{(x^2-4)^2}"


"f''(x)=\frac{(4x^3-24x)(x^2-4)^2-(x^4-12x^2)2(x^2-4)2x}{(x^2-4)^4}=
\frac{(4x^3-24x)(x^2-4)-4(x^4-12x^2)x}{(x^2-4)^3}="


"=\frac{4x^5-16x^3-24x^3+96x-4x^5+48x^3}{(x^2-4)^3}=\frac{8x^3+96x}{(x^2-4)^3}=\frac{8x(x^2+12)}{(x^2-4)^3}"





Since "f(x)" is concave up if "f''(x)>0", we conclude that "f(x)" is concave up on the region "(-2,0)\cup(2,16]."


Answer: "(-2,0)\cup(2,16]."